Mar 20, 2012 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US and Canada
    • +1-303-248-0285 : International
    • then use participant code: 3421244# (the # is needed when using the phone)
    • OR from type access code 3421244 into "join a meeting" without the # (you need java plugin)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting

  1. Beamtest
    • Draft a run plan (Elton and Zisis)

Presentations from recent meetings

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. BCAL Beam Tests 2012 - Updates since last week(Elton)
  4. Update from USM
    1. Production Manager Report (Javiera)
    2. Comments on lightguide Status (Iñaki)
    3. USM Web Interface to DB
  5. Status of Testing at JLab
    1. Hall-D ELOG
  6. Other


Attending: Elton, Tim, Beni, Yi, Eugene (JLab); Zisis, Andrei (UofR); Will, Javiera, Serguei, Hayk, Iñaki (USM)

  1. Announcements
  2. Review of Action Items
  3. BCAL Beam Tests 2012 - Updates since last week(Elton)
    1. Beam schedule in Hall B is being modified to account for loss of polarization in the new target installed last week. The plan is to run photons this week and return to electron beam next week. Schedule is fluid.
    2. Mechanical: Tom Carstens and team placed the mini-Bcal onto its support structure and practiced moves that will be required in the Hall to mount the prototype.
    3. Cables have been labeled and bundled. VXS and NIM crates have been mounted in racks in EEL 126 in preparation for cosmic testing
    4. Readout packages (SiPMs, cooling plate, preamp/summing board, temp compensation board, connectors) have been assembled. Testing is in progress.
  4. Update from USM
    1. Production Manager Report (Javiera)
      • New personnel hired and progress reported since last week
      • Ricardo has added the USM data base information to Yi's wiki page.
    2. Comments on lightguide Status (Iñaki)
      • Checking sizes of light guides
      • Found problem with dimensions
      • Start to systematic measurement of light guides on Thursday using a "profile projector" (According to Tim projects guides on a screen and allows precise measurements without touching the light guides.
      • Production of light guides with the large 5-axis CNC machine started last week. However a part failed, and will be delivered Friday.
      • The surface finish of the new machine is much better than for the 80 first articles and will save polishing time.
    3. JLab will produce tooling gauges to check overall dimensions of light guides. Will requested that we make two sets, one for USM. We will consider this request when the drawings are made.
    4. Will said that data have been taken to study systematics of the SiPM measurements and analysis is in progress
    5. USM Web Interface to DB
  5. Status of Testing at JLab
    1. Hall-D ELOG
      • Yi showed systematic studies indicating that the temperature of the SiPMs are not precisely at the temperature of the cooling plate. This was improved by reducing the gap between the sensor and the plate from 3mm to 1 mm. The silicon pad could be reduced enhancing its efficiency.
      • With the improved thermal contact the measured gain in our setup is about 7 x 10^5, considerably closer than the reported data for the 80 first article samples.
      • The impact of temperature gradients will need to be considered at the time of measurements, but also when using measurements to set conditions in the real experiment
      • It was agreed that measurements are room temperature should be less sensitive to any gradients.
      • Sergei: At lower temperatures (e.g. 5 deg) the pin contacts become problematic. They find that a few percent of MPPCs have this problem, which is a combination of sensor pins and sockets.
      • Andrei: Questioned the validity of dark rate measurements with a light source. Yi noted that the impact of dark rate on the experiment is always in the presence of signal.
  6. Other
    1. Will: Inquired about after-pulsing. They do not see any after-pulsing on the scope.
      • Yi had done a study a year ago to extract this quantity from scope pictures. He will send a link to his work, which is on the wiki. (Note added: see recent e-mail)
    2. Serguei would like to meet in a smaller group to discuss some analytic formulas used to extract quantities of interest.